OpenStudy (anonymous):

Which of the following is equal to √3√2?

OpenStudy (solomonzelman):

√3 • √2 ?

OpenStudy (solomonzelman):

(I hope not \(\sqrt{3}^{\sqrt{2}}\) )

OpenStudy (hybrik):

You can multiply √3√2 to make √6

OpenStudy (anonymous):

It looks like that, sorry xD

OpenStudy (hybrik):

You have to find the 3rd root of 2 first

OpenStudy (hybrik):

Which is 1.2599, then find the square root of that answer, which is 1.12246

OpenStudy (anonymous):

What would that be in fraction form?

OpenStudy (hybrik):

1 and 6123 over 50000

OpenStudy (anonymous):

This is what I got though?

OpenStudy (hybrik):

Oh, that would be 2 to the power 2

OpenStudy (hybrik):

Then to the cube root of that answer

OpenStudy (hybrik):

The correct way to display this is

OpenStudy (hybrik):

2 to the power of 1/6

OpenStudy (anonymous):

Ohhh! thank you, I understand where I went wrong now.

OpenStudy (hybrik):

Because you have the third root of 2, which is 2 to the power of 1/3 and then the 2nd root of 2, so 1/3*1/2 = 1/6 so, you would get 2 to the power of 1/6

OpenStudy (hybrik):

Your welcome